OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4?

OpenStudy (aroub):

y=mx+b , m=1/2 , (-2,4) 4=1/2(-2)+b 4=-1+b 5=b y=1/2x+5

OpenStudy (anonymous):

haha Thank you so much once again!